Suppose Florence has been offered a new assignment at work which comes with a $400 increase in her weekly wage. The catch is that the assignment would require her to work a bit longer so that she would need to pay for extended care for her child in the amount of $300 per week.

Further, suppose that Florence's federal marginal tax rate is 22%. She lives in Florida so she does not have to pay any state income tax. But she would also need to pay 6.2% Social Security tax and 1.45% Medicare tax on those additional earnings.

Use at least 3 decimal points in your answer and all intermediate calculations.

Given this information:

Florence's overall marginal tax rate (the sum of the above tax rates) is:___

If she accepts this new assignment she would have to pay: ___ in taxes on her additional weekly earnings.
Then her take-home after-tax additional weakly earnings would be:____.

From a purely financial standpoint, should she accept the new work assignment? Yes or No?
1
Expert's answer
2018-10-08T14:53:09-0400
Florence's overall marginal tax rate (the sum of the above tax rates) is: 22 + 6.2 + 1.45 = 29.65%.
If she accepts this new assignment she would have to pay: 400*0.2965 = $118.6 in taxes on her additional weekly earnings.
Then her take-home after-tax additional weakly earnings would be: 400 - 118.6 = $281.4.
From a purely financial standpoint, she shouldn't accept the new work assignment, because if she pay additional $300 for child care, then he will earn less than without a new assignment.